how much low fat milk (1% fat) and how much whole milk (4 % fat) should be mixed to make a 5gal of reduced fat milk (2% fat)?
x+y=5
1x+4y=5•2
Not sure if my equations are set up the correct way. I am confused on what to do since they are in percent. Please help thank you!

Found 2 solutions by rothauserc, MathTherapy:
Answer by rothauserc(4718)

Make a table
:
fat % | gallons sol'n | percent fat | total gallons fat |
1% fat | x | 0.01 | 0.01x |
4% fat | y | 0.04 | 0.04y |
mixture| x + y = 5 | 0.02 | 0.02 * 5 = 0.10 |
:
Since x + y = 5, then x = 5 - y
:
Substitute for x in the table, then
:
0.01(5 - y) + 0.04y = 0.10
:
0.05 - 0.01y + 0.04y = 0.10
:
0.03y = 0.05
:
y = 0.05/0.03 = 1.6666667 is approximately 1.67 gallon
:
x = 5 - 1.67 = 3.33 gallons
:
3.33 gallons of 1% fat mixed with 1.67 gallons of 4% fat
:

Answer by MathTherapy(10552)


how much low fat milk (1% fat) and how much whole milk (4 % fat) should be mixed to make a 5gal of reduced fat milk (2% fat)?
x+y=5
1x+4y=5•2
Not sure if my equations are set up the correct way. I am confused on what to do since they are in percent. Please help thank you!
Answer to  your QUESTION: "Yes, your equations are correct!" What you did was multiply the percents by 100, thus making 1%, 1, 4%, 4, and 2%, 2.
Now just continue to solve for each type of milk.
You can either use ELIMINATION or SUBSTITUTION. You choose!!